OpenStudy (anonymous):

1.cot x sec^4 x = cot x + 2 tan x + tan^3 x

OpenStudy (anonymous):

The cots cancel

OpenStudy (anonymous):

sec^4 x = 2 tan x + tan^3 x

OpenStudy (anonymous):

tan x = sin x / cos x

OpenStudy (anonymous):

So \[(\frac{ 1 }{ \cos x})^4 = 2 \frac{ \sin x }{ \cos x } + (\frac{ \sin x }{ \cos x })^3\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Not sure where to go from here

OpenStudy (loser66):

What are we supposed to do? prove? One more thing: cannot cancel out cot on that way because from the left hand side it is *, and from the right hand side it is +

OpenStudy (loser66):